A 15 L container is filled with helium and the pressure is 154 atm and the temperature is 311 K. What volume can be filled in liters when the temperature is 25° C and the atmospheric pressure is 759 torr?



1
Expert's answer
2021-03-02T02:00:39-0500

atm = torr / 760

T(K) = T(°C) + 273.15


P1 = 154 atm

V1 = 15 L

T1 = 311 K

P2 = 759 torr = 0.9987 atm

V2 = unknown

T2 = 25°C = 298 K

 

Solution:

The Combined Gas Law:

P1V1/T1 = P2V2/T2

Cross-multiply to clear the fractions:

P1V1T2 = P2V2T1

Divide to isolate V2:

V2 = (P1V1T2) / (P2T1)

Plug in the numbers and solve for V2:

V2 = (154 atm × 15 L × 298 K) / (0.9987 atm × 311 K) = 2216.3 L

V2 = 2216.3 L


Answer: The volume is 2216.3 liters.